7) Which of the following is an example of value-added work?
A) Repairing a defective product B) Unpacking boxes in a stock room C) Waiting for customers to order D) Taking orders from customers
 
8) Which of the following can NOT be used to determine the extent to which an operation is lean?
A) Overall equipment effectiveness B) Value-added percentage C) Inventory turns D) Labor content
 
9) A product test is designed in such a way that for a defective product to be undiscovered, all four inspections would have to fail to catch the defect. The probability of catching the defect in inspection I is 90%, in inspection 2 80%, in inspection 3, 70% and in inspection 4.90%. What is the probability of catching a defect? Show work for partial credit.
 
10) Which of the following information is NOT included in Gantt charts?
A) The expected completion time of the project B) The critical path C) The slack time D) The budget for the project
 
11) Which of the following is the Pareto principle?
A) The majority of problems can be explained by many symptoms. B) The majority of problems can be explained by a few root causes. C) The majority of problems can be explained by many root causes. D) The majority of problems can be explained by one root cause.
 
12) C&A uses an assembly line of 30 stations to make a product. The cycle time is 5 minutes per unit. Step 10 is the most error-prone operation. Inspection occurs at the last station. What will be the information turnaround time for a defect made at station 10 if there is no inventory between stations? Show work for partial credit
Answer #1

7)

Correct answer is D) Taking orders from customers

Taking orders from customers is a value-added work

8)

correct answer is D) Labor content

Labor content can NOT be used to determine the extent to which an operation is lea

9)

probability of catching a defect = 0.9*0.8*0.7*0.9 = 0.4536

10)

Correct answer is A) the expected completion time of the project

the expected completion time of the project is not shown in Gantt chart

11)

Correct answer is B) The majority of problems can be explained by a few root causes

Pareto chart says 80% of issues are created by 20% root causes

12)

There are 20 stations between station 10 and last station as cycle time is 5 minutes so information turnaround time = 20*5 = 100 minutes